# Binary Flow Matching: Prediction-Loss Space Alignment for Robust Learning

### v1

@@ I Introduction
Flow matching (FM) [11, 13] and diffusion models [7, 20] provide a unified framework for transforming simple noise distributions into complex data manifolds through continuous probability paths. While theoretically applicable to arbitrary distributions, their most prominent empirical successes have relied on Gaussianization pre-processing, where linear interpolation facilitates smooth transport trajectories. Extending these continuous frameworks to discrete or binary data remains an open challenge. Existing approaches largely fall into two categories: those that explicitly redesign the forward process for discrete state spaces via transition kernels [1, 21, 2], and those that treat binary data as ”analog bits” [3], embedding them into continuous Euclidean space. While the latter preserves architectural simplicity by leveraging standard Gaussian interpolation, its optimal objective design and numerical stability remain poorly understood. Recently, the ”back-to-basics” philosophy exemplified by the Just Image Transformer (JiT) [10] demonstrated that for continuous signals, a remarkably simple design—signal-space prediction ($x$-prediction)—achieves state-of-the-art performance. A critical question naturally arises: Does the inherent advantage of $x$-prediction still exist under binary distributions, and is the paradigm fundamentally robust?
@@
Second, we identify a foundational robustness issue: pairing $x$-prediction with the prevailing velocity-matching loss ($v$-loss) induces a prediction–loss space mismatch. We rigorously demonstrate that this mismatch leads to a structural singularity where gradient variance exhibits a third-order divergence as the flow approaches the target. We reveal that this pathology is a general characteristic of the $x$-prediction paradigm, and its stability in current literature often relies on boundary-avoiding heuristics, such as the Logit-Normal sampling strategies used in JiT [10] and Stable Diffusion 3 [6].
Third, we establish the principles for robust and topology-aware learning on binary domains. We propose prediction-loss space alignment—shifting the objective back to the signal space—as the structural solution that ensures stability independent of the sampling distribution. Furthermore, we show that once alignment is secured, the optimal loss design is dictated by the signal topology: geometric regression (Mean Squared Error, MSE) is uniquely suited for recovering spatially correlated structures like binary images, whereas probabilistic objectives (e.g., Binary Cross-Entropy, BCE) provide superior inductive biases for distinguishing independent symbolic states in tasks such as Multiple-Input-Multiple-Output (MIMO) detection.
@@
  • Validation of $x$-Prediction for Binary Data: We provide an empirical validation that signal-space prediction ($x$-prediction) is a superior paradigm for binary manifolds. We show that directly parameterizing the model in signal space offers improved generative quality and path consistency compared to traditional velocity prediction for discrete-latent data.
  • Diagnosis of Mismatch Singularity: We identify a foundational robustness issue: pairing $x$-prediction with the prevailing velocity-matching loss ($v$-loss) induces a prediction-loss space mismatch. We rigorously demonstrate that this mismatch leads to a structural singularity in the gradient variance. This pathology is a general characteristic of the $x$-prediction paradigm, and we reveal that its stability in current literature often relies on boundary-avoiding heuristics, such as Logit-Normal sampling or timestep clipping.
  • Principles of Aligned and Topology-Aware Learning: We establish the principles for robust learning on binary domains. We propose prediction-Loss space alignment—shifting the objective back to the signal space—as the structural solution that ensures stability independent of the sampling distribution. Furthermore, we show that once alignment is secured, the optimal loss design is dictated by the signal topology: probabilistic objectives (e.g., BCE) are suited for independent symbolic recovery, whereas geometric regression (MSE) more effectively preserves spatial correlations.
@@ II-A Continuous Diffusion and Flow Matching
However, our work identifies a critical structural limit to this specific $x$-prediction paradigm. We demonstrate that while this combination is effective for continuous manifolds, the velocity-based loss becomes pathologically singular when applied to binary priors under $x$-prediction. We note that recent high-performance models like Stable Diffusion 3 [6] and JiT [10] both rely on a Logit-Normal sampling trick to bias training towards intermediate timesteps ($t\approx 0.5$). While these works use such samplers to improve empirical results or hide boundary instabilities, our work provides a theoretical refinement of the $x$-prediction framework. We prove that by implementing prediction-loss space alignment—shifting the loss calculation back to the signal space—we can analytically cancel the singularity. This refinement allows $x$-prediction models to achieve superior stability and sampling quality on binary domains without requiring specialized sampling schedules, providing a more principled foundation for robust learning.
@@ II-B Discrete and Binary Diffusion Models
To bridge continuous diffusion with discrete data, early research focused on categorical transition kernels [1, 14]. A simpler alternative, “Analog Bits” [3], treats binary data as continuous values, enabling the use of standard Gaussian paths. Within this “Analog” framework, the choice of loss remains a point of contention. Bit Diffusion [3] utilizes MSE for binary distributions, while CDCD [4] advocates for Cross-Entropy to ensure stable gradients in categorical settings like text.
Our work unifies these diverse strategies by identifying signal topology as the key determinant for loss selection. We show that once Prediction-Loss Alignment is secured, the choice between BCE and MSE is dictated by the data’s internal correlation structure. We demonstrate that $x$-space prediction is inherently superior for binary settings because it allows for direct probabilistic interpretation (BCE), and our alignment framework ensures that this process remains robust throughout the entire training trajectory.
@@ IV-A Preliminaries
Assumption IV.3 (Asymptotic Residual Scaling).
The decay rate of the minimum residual $R(t)=\mathbb{E}[\|\hat{\mathbf{x}}_{\theta}-\mathbf{x}\|^{2}]$ as $t\to 1$ is dictated by the signal’s probability support:
@@
    Continuous Correlated Case: For standardized signals $\mathbf{x}\sim\mathcal{N}(0,\mathbf{\Sigma})$, the Bayes-optimal residual vanishes quadratically: $R(t)=\mathcal{O}((1-t)^{2})$.
@@
    Binary Discrete Case: For binary data $\mathbf{x}\in\{-1,1\}^{D}$, any finite-Lipschitz network $\hat{\mathbf{x}}_{\theta}$ maintains a non-vanishing residual $\epsilon^{2}>0$ as $t\to 1$: $R(t)=\Omega(1)$.
@@ IV-B Divergence Analysis: The Mismatch Singularity
    For binary signals, $\mathcal{I}$ exhibits a third-order divergence ($\mathcal{O}((1-t)^{-3})$).
Proof Sketch. The integral $\mathcal{I}\propto\int_{0}^{1}(1-t)^{-4}R(t)dt$ is determined by the asymptotic behavior of the residual $R(t)$ as $t\to 1$. For continuous signals, $R(t)\sim(1-t)^{2}$ partially regularizes the integrand to $(1-t)^{-2}$. For binary data, the persistent approximation error $R(t)\sim\Omega(1)$ leaves the $(1-t)^{-4}$ term uncompensated, triggering explosive divergence. In both cases, the configuration is structurally ill-conditioned. (See Appendix A for full proof).
@@ IV-C Implicit Stabilization via Logit-Normal Sampling
JiT’s empirical successes in $x$-prediction with $v$-loss parameterization [10] rely on the Logit-Normal sampling schedule, originally proposed in [6] for importance sampling based on signal-to-noise ratios. A Logit-Normal density $\pi_{LN}(t;m,s)$ is defined as:
$\begin{split}\pi_{LN}(t)&=\frac{1}{s\sqrt{2\pi}}\frac{1}{t(1-t)}\exp\left(-\frac{(u-m)^{2}}{2s^{2}}\right),\\ u&=\ln\frac{t}{1-t},\quad u\in\mathbb{R}.\end{split}$ (4)
We reveal that while its primary intent is to reweight timesteps by learning difficulty, it accidentally acts as a numerical ”safety valve” for the mismatched objective.
@@
Let $\pi_{LN}(t)$ denote a Logit-Normal sampling density over $t\in(0,1)$ with scale parameter $s>0$ and mean $m=0$, i.e.,
$t=\sigma(u),\quad u\sim\mathcal{N}(m,s^{2}),\quad\sigma(u)=\frac{1}{1+e^{-u}}.$ (5)
This distribution provides a structural suppression of the boundary $t\to 1$, rendering the weighted variance integral
@@
    Binary Case: For discrete binary data, the residual remains constant $R(u)\sim\Omega(1)$, yielding an integrand proportional to $\exp(-u^{2}/(2s^{2})+4u)$.
The peak of the effective density occurs at $u_{\mathrm{peak}}=ns^{2}$ with $n=2$ for continuous and $n=4$ for binary data. Thus, the Logit-Normal scale $s$ determines how far into the boundary region the sampling density remains significant. Convergence of $\mathcal{I}_{\pi}$ is guaranteed for any finite $s>0$, but the practical sampling budget near $t\approx 1$ is reduced for binary data due to the stronger third-order singularity.
@@ IV-D Stability via Prediction-Loss Space Alignment
Consider an aligned training configuration where the objective is defined in the network’s prediction space. Under Assumption IV.1, the stochastic gradient $\mathbf{g}_{t}$ is uniformly bounded, ensuring sampler-agnostic stability:
@@
    Binary Manifolds: Signal alignment using either MSE or BCE objectives yields $\mathbb{E}[\|\mathbf{g}_{t}\|^{2}]=\mathcal{O}(1)$, effectively eliminating the third-order divergence.
Proof Sketch. Alignment ensures the algebraic factor $(1-t)^{-2}$ is absent from the gradient $\mathbf{g}_{t}$. In continuous domains, this restores stability to the $x$-prediction paradigm popularized by JiT. In binary domains, alignment prevents the non-vanishing residual $R(t)\sim\Omega(1)$ from triggering the mismatch singularity, allowing for robust convergence under uniform sampling. (See Appendix C for full proof).
Figure 2: Toy experiments reveal that Logit-Normal sampling stabilizes training by suppressing boundary singularities. (a,b) Gradient norms under Logit-Normal vs. uniform sampling for binary and Gaussian data, respectively. Uniform sampling leads to severe gradient explosion under $x$-prediction with velocity loss, while Logit-Normal sampling yields bounded gradients. (c,d) Corresponding training loss trajectories show that Logit-Normal sampling enables stable convergence, whereas uniform sampling results in highly unstable optimization. (e) Distribution of sampled $t$ values under Logit-Normal and uniform schedules, illustrating strong suppression of the boundary region $t\to 1$. (f,g) Gradient norms under uniform sampling for different prediction–loss pairings, confirming that instability is specific to mismatched $x$-prediction with velocity loss. (h,i) Training losses under uniform sampling further highlight that aligned objectives remain stable. (j) Bit error rate (BER) as a function of $t$ in the binary case, compared with the Bayes-optimal MMSE estimator, showing that performance degradation concentrates near the singular boundary region. Note that all prediction-loss-aligned approaches achieve MMSE except $x$-prediction with $v$-loss.
@@ Discussion.
Once prediction–loss alignment guarantees stable optimization, the loss function selection is dictated by the intrinsic topology of the binary signal. Probabilistic objectives such as BCE are naturally suited for independent symbolic recovery, whereas geometric regression losses such as MSE are more appropriate when binary data exhibit structured or correlated organization.
V Toy Experiments: Controlled Verification of Sampling-Induced Stability
V-A Problem Setup
We first conduct controlled toy experiments to verify the theoretical predictions on gradient instability and sampling-induced stabilization. Rather than evaluating generation quality, these experiments focus on optimization dynamics, in particular the behavior of gradient norms and training losses under different parameterization–loss pairings and $t$-sampling strategies.
We consider two synthetic data manifolds: continuous Gaussian vectors and binary $\{-1,+1\}^{D}$ vectors. In both cases, we train a lightweight conditional MLP to perform $x$-prediction under velocity-based loss, which corresponds to the mismatched configuration analyzed in Theorem IV.4. We compare uniform sampling of the time variable $t$ with Logit-Normal sampling, which suppresses the boundary region $t\to 1$. (See Appendix D for detailed settings.)
Loss $v$-MSE (uniform) BCE (w) CE (uniform) $x$-MSE $v$-MSE
Pred / Sample $x$-pred / uniform Logit-Normal $x$-pred / uniform $x$-pred / uniform $v$-pred / uniform
FID $\downarrow$ 1247.9823 5.5797 835.7452 991.7779 4.9386 8.7199
TABLE I: FID on Binary MNIST (lower is better). All results use the same training schedule and sampling procedure, and use the best checkpoint (lowest validation loss).
Figure 3: Binary MNIST qualitative samples and training dynamics across different objectives. Each column illustrates a specific parameterization–loss pairing: (1) Mismatched (Uniform): $x$-prediction with $v_{\mathrm{MSE}}$ loss, exhibiting immediate catastrophic divergence; (2) Mismatched (Logit-Normal): while the Logit-Normal schedule masks the singularity during training (blue), the validation loss (red) exhibits extreme, orders-of-magnitude oscillations, revealing an ill-conditioned vector field; (3–4) Aligned (BCE): stable convergence but yields thick strokes due to the independent Bernoulli assumption; (5) Aligned (MSE): our proposed method, achieving the lowest FID and most stable convergence; (6) Original FM: standard $v$-prediction with $v_{\mathrm{MSE}}$ loss. The sharp contrast between the instability in (2) and the monotonic convergence in (5) empirically validates that Alignment is the fundamental remedy for the structural singularity, proving that boundary-avoiding heuristics alone cannot resolve the underlying numerical fragility.
V-B Results and Discussion

### v2

@@ I Introduction
Flow matching (FM) [12, 14] and diffusion models [8, 21] provide a unified framework for transforming simple noise distributions into complex data manifolds through continuous probability paths. While theoretically applicable to arbitrary distributions, their most prominent empirical successes have relied on Gaussianization pre-processing, where linear interpolation facilitates smooth transport trajectories. Extending these continuous frameworks to discrete or binary data remains an open challenge. Existing approaches largely fall into two categories: those that explicitly redesign the forward process for discrete state spaces via transition kernels [1, 22, 2], and those that treat discrete data as “analog bits” [3, 6], embedding them into continuous Euclidean space. While this line preserves architectural simplicity and has already produced meaningful progress (e.g., CatFlow [6]), how different objective choices affect generation quality under different discrete data topologies remains underexplored. Recently, the ”back-to-basics” philosophy exemplified by the Just Image Transformer (JiT) [11] demonstrated that for continuous signals, a remarkably simple design—signal-space prediction ($x$-prediction)—achieves state-of-the-art performance. A critical question naturally arises: Does the inherent advantage of $x$-prediction still exist under binary distributions, and is the paradigm fundamentally robust?
@@
Second, we identify a foundational robustness issue: pairing $x$-prediction with the prevailing velocity-matching loss ($v$-loss) induces a mismatch between the prediction and loss spaces. We rigorously demonstrate that this mismatch creates a structural singularity whose severity depends on terminal residual scaling during training (reaching third-order divergence in the strong binary regime). This analysis helps explain why training can become sensitive near the terminal region, and why boundary-avoiding heuristics (e.g., Logit-Normal timestep sampling as used in JiT [11] and Stable Diffusion 3 [7]) are often beneficial in practice.
Third, we establish the principles for robust and topology-aware learning on binary domains. We propose prediction-loss space alignment—shifting the objective back to the signal space—as the structural solution that ensures stability independent of the sampling distribution. Furthermore, we show through theory and experiments that once alignment is secured, the optimal loss design is dictated by the signal topology: geometric regression (Mean Squared Error, MSE) is uniquely suited for recovering spatially correlated structures like binary images, whereas probabilistic objectives (e.g., Binary Cross-Entropy, BCE) provide superior inductive biases for distinguishing independent symbolic states in tasks such as Multiple-Input-Multiple-Output (MIMO) detection.
@@
  • $x$-Prediction on Binary Data. We show that signal-space prediction remains effective on binary manifolds, outperforming velocity prediction in generative quality and path consistency.
  • Mismatch Analysis. We show that coupling $x$-prediction with velocity matching induces a structural singularity, clarifying why boundary-avoiding timestep sampling can stabilize training.
  • Alignment and Loss Design. We propose prediction-loss space alignment as a sampler-agnostic remedy, and show that aligned loss choice should follow signal topology: BCE for independent symbolic recovery, MSE for correlated binary manifolds.
@@ II-A Continuous Diffusion and Flow Matching
Our work complements this line by analyzing a setting where $x$-prediction is transferred to binary priors. We show that while the combination is effective, velocity-based supervision under $x$-prediction introduces a terminal-time singular weighting. We further provide a theoretical refinement via prediction-loss alignment—shifting the loss calculation back to the signal space—which analytically removes this singular term. This yields stable training on binary as well as continous-near-gaussian domains under uniform timestep sampling and offers a principled explanation for why boundary-biased schedules (e.g., Logit-Normal) can help in practice.
@@ II-B Discrete and Binary Diffusion Models
To bridge continuous diffusion with discrete data, early research focused on categorical transition kernels [1, 15]. A complementary alternative, the “Analog Bits” line [3, 6], treats binary or discrete data as continuous values, enabling standard Gaussian paths and strong empirical performance. In particular, VFM [6] provides an important variational construction that supports practical $x$-prediction with cross-entropy-style objectives, while also encompassing the original flow-matching setting ($v$-prediction with $v$-loss) in a unified formulation. Within this broader analog/variational line, different objectives have been explored for different domains: Bit Diffusion [3] uses MSE for binary image-style settings, while CatFlow/VFM [6] and CDCD [4] adopts Cross-Entropy-style supervision in categorical settings such as text.
Our work is complementary to these strategies. Rather than proposing another path construction, we explicitly characterize how prediction-loss mismatch affects optimization stability, and show that prediction-loss alignment removes the singular weighting. Building on the unified objective perspective above, we further provide both theoretical and empirical evidence that the BCE-vs.-MSE choice should be determined by data topology (independent symbolic states vs. spatially correlated structures), yielding a practical criterion for loss selection in binary settings.
@@ IV-A Preliminaries
Assumption IV.3 (Training-Stage Terminal Residual Regimes).
Let
$R_{\theta}(t):=\mathbb{E}\!\left[\|\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)-\mathbf{x}\|^{2}\right]$
denote the training-stage residual attained by the current network parameters $\theta$. Its terminal scaling as $t\to 1$ follows practically relevant regimes:
@@
    Continuous Correlated Case (Bayes-optimal benchmark): For standardized signals $\mathbf{x}\sim\mathcal{N}(0,\mathbf{\Sigma})$, the Bayes-optimal residual satisfies $R^{\star}(t)=\mathcal{O}((1-t)^{2})$ near the terminal region.
@@
    Binary Regime A (no end-to-end residual skip): In early training, if the architecture does not provide an explicit end-to-end residual path from input $\mathbf{z}_{t}$ to output, then no identity-like projection is available near initialization, and $R_{\theta}(t)\gtrsim c_{0}>0$ on a terminal interval.
  3. 3.
    Binary Regime B (with end-to-end residual skip): If the architecture includes an explicit end-to-end residual path (e.g., identity skip plus zero-initialized residual branch), then near initialization $\hat{\mathbf{x}}_{\theta}\approx\mathbf{z}_{t}$ (identity-like projection), and $R_{\theta}(t)=\mathcal{O}((1-t)^{2})$.
@@ IV-B Divergence Analysis: The Mismatch Singularity
    For binary signals in Regime A, $\mathcal{I}$ exhibits a third-order divergence ($\mathcal{O}((1-t)^{-3})$).
  3. 3.
    For binary signals in Regime B, $\mathcal{I}$ exhibits a first-order divergence ($\mathcal{O}((1-t)^{-1})$).
Proof Sketch. The mismatch gradient second moment satisfies
$\mathbb{E}\|\mathbf{g}_{t}\|^{2}\gtrsim(1-t)^{-4}R_{\theta}(t),$
so $\mathcal{I}\propto\int_{0}^{1}(1-t)^{-4}R(t)\,dt$ is governed by terminal residual scaling. For continuous signals, even under the Bayes-optimal benchmark $R^{\star}(t)\sim(1-t)^{2}$, the integrand scales as $(1-t)^{-2}$ and $\mathcal{I}$ still diverges at first order. For binary Regime B, $R_{\theta}(t)\sim(1-t)^{2}$ gives the same first-order divergence, while binary Regime A with $R_{\theta}(t)\gtrsim c_{0}$ yields third-order divergence. Hence, mismatch-induced singular amplification is structural, while severity depends on residual scaling. (See Appendix A for full proof).
@@ IV-C Implicit Stabilization via Logit-Normal Sampling
JiT’s empirical successes in $x$-prediction with $v$-loss parameterization [11] rely on the Logit-Normal sampling schedule, originally proposed in [7] for importance sampling based on signal-to-noise ratios. Let
$u\sim\mathcal{N}(m,s^{2}),\qquad t=\sigma(u)=\frac{1}{1+e^{-u}}.$ (4)
Then $t\in(0,1)$ follows a Logit-Normal distribution with density
$\begin{split}\pi_{LN}(t;m,s)&=\frac{1}{s\sqrt{2\pi}}\frac{1}{t(1-t)}\exp\!\left(-\frac{(\operatorname{logit}(t)-m)^{2}}{2s^{2}}\right),\\ \operatorname{logit}(t)&=\ln\frac{t}{1-t}.\end{split}$ (5)
We reveal that while its primary intent is to reweight timesteps by learning difficulty, it accidentally acts as a numerical “safety valve” for the mismatched objective.
@@
Let $\pi_{LN}(t;0,s)$ be the Logit-Normal density defined above on $t\in(0,1)$. This distribution provides a structural suppression of the boundary $t\to 1$, rendering the weighted variance integral
@@
    Binary Case: In Regime A, $R_{\theta}(u)\sim\Omega(1)$, yielding $\exp(-u^{2}/(2s^{2})+4u)$; in Regime B, $R_{\theta}(u)\sim e^{-2u}$, yielding $\exp(-u^{2}/(2s^{2})+2u)$.
The peak of the effective density occurs at $u_{\mathrm{peak}}=ns^{2}$, with $n=2$ for continuous data (and binary Regime B), and $n=4$ for binary Regime A. Thus, the Logit-Normal scale $s$ determines how far into the boundary region the sampling density remains significant. Convergence of $\mathcal{I}_{\pi}$ is guaranteed for any finite $s>0$, while the practical sampling budget near $t\approx 1$ is most constrained in binary Regime A.
@@ IV-D Stability via Prediction-Loss Space Alignment
Consider an aligned training configuration where the objective is defined in the network’s prediction space. Under Assumption IV.1, the stochastic gradient $\mathbf{g}_{t}$ is uniformly bounded, ensuring sampler-agnostic stability (independent of the residual-scaling regime in Assumption IV.3):
@@
    Binary Manifolds (Regime A/B): Signal alignment using either MSE or BCE objectives yields $\mathbb{E}[\|\mathbf{g}_{t}\|^{2}]=\mathcal{O}(1)$ in both Regime A and Regime B, effectively eliminating mismatch-induced divergence.
Proof Sketch. Alignment ensures the algebraic factor $(1-t)^{-2}$ is absent from the gradient $\mathbf{g}_{t}$. In continuous domains, this restores stability to the $x$-prediction paradigm popularized by JiT. In binary domains, this removes mismatch-induced singular amplification irrespective of whether training is in Regime A or Regime B, allowing robust convergence under uniform sampling. (See Appendix C for full proof).
@@ Discussion.
Once prediction–loss alignment guarantees stable optimization, the loss function selection is dictated by the intrinsic topology of the binary signal. This can be interpreted from a Bayesian posterior perspective under the Gaussian corruption path
$\mathbf{z}_{t}=t\mathbf{x}+(1-t)\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),$
which gives
$p(\mathbf{x}\mid\mathbf{z}_{t})\propto p(\mathbf{x})\exp\!\left(-\frac{\|\mathbf{z}_{t}-t\mathbf{x}\|^{2}}{2(1-t)^{2}}\right).$
If the prior is approximately factorized, $p(\mathbf{x})\approx\prod_{i}p(x_{i})$, then a coordinate-wise Bernoulli posterior model is appropriate, favoring BCE. If dimensions are strongly coupled so that $p(\mathbf{x}\mid\mathbf{z}_{t})\neq\prod_{i}p(x_{i}\mid\mathbf{z}_{t})$, the factorized Bernoulli assumption is misspecified, and MSE is often the better aligned objective for structured denoising. Therefore, probabilistic objectives such as BCE are naturally suited for independent symbolic recovery, whereas geometric regression losses such as MSE are more appropriate when binary data exhibit structured or correlated organization.
IV-F Controlled Verification on Toy Problems
We use toy problems only to isolate the instability mechanism, not to provide a comprehensive comparison of downstream binary objectives. We use Gaussian and binary synthetic manifolds with a lightweight conditional MLP under the standard interpolation path; full architectures, losses, and evaluation details are deferred to Appendix D-A. The Gaussian toy isolates the effect of timestep sampling on the mismatched $x$-prediction+$v$-MSE setup, whereas the binary toy fixes uniform sampling and compares prediction–loss pairings.